Which property justifies the following statement.

If 5a-10b=32 then 5(a-2b)=32

A. Commutative Property

B. Associative Property

C. Distributive Property

D. Inverse Property

thank you

Answer by RAY100(1637)

The distribution Property is usually stated as a(c+b) = ac +bc
.
the reverse of this property is factoring , which is what we did, taking the 5 out.
.
The other properties are cummulative, a*b = b*a
associative (a*b)*c = a*(b*c)
inverse ,,reciprocal of a = 1/a